OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form and show your work to receive full credit. 1) 2x^4(4x^2 + 3x + 1) 2) (4x – 3)(2x^2 – 7x + 1) 3) (x^2 + 4x – 3)(2x^2 + x + 6)

OpenStudy (anonymous):

That's wrong lol.

OpenStudy (lgbasallote):

@rebeccaskell94

OpenStudy (lgbasallote):

only way to simplify #1 is to factor out that quadratic expression in the parenthesis

OpenStudy (anonymous):

what do you mean

OpenStudy (lgbasallote):

factor out 4x^2 + 3x + 1

OpenStudy (anonymous):

7x^2 + 1?

OpenStudy (lgbasallote):

did you just add? that's not how it works

OpenStudy (lgbasallote):

you cant add x^2 and x

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

Multiply?

OpenStudy (lgbasallote):

wait...i'll show you why you can't add...okay....let me show you why...let's say we have x^2 + x ....x can be any number right?

OpenStudy (anonymous):

I'd say 1, but yeah.

OpenStudy (anonymous):

Then it would be x^3

OpenStudy (lgbasallote):

what do you mean 1?

OpenStudy (anonymous):

I'd say x as 1.

OpenStudy (lgbasallote):

hmm 1 really cant be used for proving...so we'll use 2...according to your theory it will be x^3...let's say x is 2 so x^3 = 2^3 = 2x 2 x2= 8 BUT x^2 + x is equal to 2^2 + 2 = 4 + 2 = 6...see why you cant add?

OpenStudy (lgbasallote):

anyway on to factoring now that that's cleared

OpenStudy (lgbasallote):

4x^2 + 3x + 1 that is also equal to 4x^2 + 2x + x + 1 right??

OpenStudy (anonymous):

no..

OpenStudy (lgbasallote):

no? why not?

OpenStudy (anonymous):

Why would you put x by it's self from 3x??

OpenStudy (lgbasallote):

3x = 2x + x

OpenStudy (anonymous):

Like, 4x^2 + 3x + 1. Why would you take 1x from 2x put the x by it's self and have just an x?

OpenStudy (anonymous):

I'm so confused.q

OpenStudy (anonymous):

but it has 3x. . . why say 3x is 2x + x..

OpenStudy (lgbasallote):

because 3 = 2 + 1

OpenStudy (lgbasallote):

3x = 2x + 1x 3 = 2 + 1 seeing it?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

2x4(4x2 + 3x + 1)

OpenStudy (lgbasallote):

so you're good with that now? 3x = 2x + x?

OpenStudy (anonymous):

2x^ 4(4x^2 + 3x + 1)

OpenStudy (anonymous):

Ugh yeah I am

OpenStudy (lgbasallote):

4x^2 + 2x + x + 1 <---that's equal to 4x^2 + 3x + 1

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

now we "factor out" 2x from 4x^2 + 2x 4x^2 + 2x = 2x(2x +1) agree?

OpenStudy (anonymous):

But it doesn't ask me to factor them out, it just asks me to put them in standard form and show my work..

OpenStudy (lgbasallote):

it says simplify o.O

OpenStudy (anonymous):

Oh. okay..

OpenStudy (lgbasallote):

if it's just standard form then what the answerer said about 8x^6 + 6x^5 + 2x^4 is right

OpenStudy (lgbasallote):

this question is weird o.O asks to simplify at the same time put in standard form @_@

OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form and show your work to receive full credit.

OpenStudy (anonymous):

The FINAL product in standard form...

OpenStudy (lgbasallote):

im sorry my brain is fried today...lemme get you someone else @_@ @ParthKohli please help

OpenStudy (anonymous):

:[ okay.

Parth (parthkohli):

See, you basically do two things here: 1) You can first multiply the term outside the bracket with the terms inside. 2) You can first do quadratic grouping and keep the term outside the bracket.

OpenStudy (anonymous):

I don't get it..